# POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization

### v2

@@ Abstract
Online polarization poses a growing challenge for democratic discourse, yet most computational social science research remains monolingual, culturally narrow, or event-specific. We introduce POLAR, a multilingual, multicultural, and multievent dataset with over 23k instances in seven languages from diverse online platforms and real-world events. Polarization is annotated along three axes: presence, type, and manifestation, using a variety of annotation platforms adapted to each cultural context. We conduct two main experiments: (1) we fine-tune six multilingual pretrained language models in both monolingual and cross-lingual setups; and (2) we evaluate a range of open and closed large language models (LLMs) in few-shot and zero-shot scenarios. Results show that while most models perform well on binary polarization detection, they achieve substantially lower scores when predicting polarization types and manifestations. These findings highlight the complex, highly contextual nature of polarization and the need for robust, adaptable approaches in NLP and computational social science. All resources will be released to support further research and effective mitigation of digital polarization globally.
POLAR: A Benchmark for Multilingual, Multicultural, and Multi-Event Online Polarization
Usman Naseem1, Juan Ren1, Saba Anwar2, Sarah Kohail6, Rudy Alexandro Garrido Veliz2, Robert Geislinger2, Aisha Jabr6, Idris Abdulmumin5, Laiba Qureshi2, Aarushi Ajay Borkar2, Maryam Ibrahim Mukhtar7, Abinew Ali Ayele2,3, Ibrahim Said Ahmad7,8, Adem Ali2,3, Martin Semmann2, Shamsuddeen Hassan Muhammad4,7, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Bahir Dar University, 4Imperial College London, 5University of Pretoria, 6Zayed University, 7Bayero University Kano, 8Northeastern University
@@ 1 Introduction
Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (waller2021quantifying; iandoli2021impact). It is often fueled by biased or inflammatory content on social media, reinforcing echo chambers and undermining mutual understanding (garimella2018polarization). Polarized discourse not only amplifies ideological divides but can also escalate into hate speech, harassment, and real-world violence. As such, early detection of polarization is critical to designing interventions that promote healthier online ecosystems.
@@
Figure 1: Pipeline for POLAR construction: multi-platform data curation in 7 languages, annotation workflow with quality control, and benchmarking.
@@
Despite growing attention, computational approaches to polarization suffer from major limitations. First, most existing datasets focus on English or high-resource languages, reflecting a widespread trend across NLP tasks that ignores the rich diversity of linguistic and sociocultural contexts in which polarization manifests. Second, current benchmarks are often event-specific or monodomain, such as U.S. elections or Western political debates, limiting their generalizability. Third, the conceptualization of polarization in NLP has largely been binary or topic-focused, overlooking the multifaceted ways in which polarization is expressed through vilification, dehumanization, stereotyping, or other rhetorical tactics.
To address these gaps, we introduce POLAR a novel multilingual, multicultural, and multievent dataset for fine-grained polarization detection. It spans seven languages across diverse regions, including low-resource languages such as Amharic and Hausa. Our data is sourced from various platforms (e.g., Twitter/X, Facebook, BlueSky, Reddit, and local news outlets), reflecting authentic, event-driven discourse ranging from armed conflict (e.g., the Tigray War) to social justice movements (e.g., abortion rights, migration crises).
Unlike prior work, POLAR supports three complementary tasks:
  1. 1.
    Binary Polarization Detection: Is a message polarized or not?
  2. 2.
@@
    Polarization Type Classification: What social dimension is targeted (e.g., political, religious, racial)?
  3. 3.
@@
    Manifestation Identification: How is polarization rhetorically expressed (e.g., stereotyping/deindividuation, vilification, dehumanization, extreme language)?
@@
We design a cross-cultural annotation protocol tailored for each language’s sociopolitical context. The complete pipeline can be found in Figure 1. We benchmark a range of multilingual language models (MLMs) and large language models (LLMs) under zero-shot, few-shot, and cross-lingual scenarios. Our experiments highlight the challenges of generalization and the limitations of current models in capturing nuanced rhetorical patterns across languages. Our contributions are as follows:
@@
  • We release POLAR, the first large-scale, multilingual, fine-grained dataset for polarization detection across 7 languages and diverse global events.
@@
  • We provide comprehensive benchmarks using state-of-the-art MLMs and LLMs across multiple evaluation settings (monolingual, cross-lingual, and few-shot).
Figure 2: Languages represented in the dataset, covering diverse linguistic and regional contexts.
@@ 2 Related Work
Online polarization has long been recognized as a threat to democracy and social cohesion, intensifying through social media echo chambers and biased content (waller2021quantifying; iandoli2021impact; garimella2018polarization). As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly. Such efforts are critical not only for content moderation, but also for peacebuilding, policy development, and responsible digital governance. Foundational research has defined polarization as both intergroup hostility and ingroup cohesion (arora2022polarization), and has highlighted its relationship with hate speech, fragmentation, and incivility (Mathew2021HateXplain).
A growing body of research has documented the role of online spaces in intensifying polarization across different regions (kubin2021role; barbera2020social; gitlin2016outrage; soares2021hashtag). However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (kubin2021role). This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially in the Global South or multilingual regions.
The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities. Recent shared tasks on hate speech and toxicity (basile2019semeval; pamungkas2020misogynistic; ousidhoum2024semeval) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations. Our work addresses this gap by presenting the first comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
@@ 3 POLAR Dataset Construction
We aim to construct a multilingual, multi-event, and multi-platform social media dataset spanning multiple continents. The dataset is designed to cover Indo-European and Afro-Asiatic language families, with a particular focus on balancing linguistic diversity and resource availability. It includes high-resource languages such as English, Spanish, and German; medium-resource languages like Arabic and Urdu; and low-resource languages such as Hausa and Amharic. While several Indo-European languages are represented (e.g., English, Spanish, German, Urdu), the dataset also incorporates key Afro-Asiatic languages (e.g., Hausa, Arabic, Amharic) to ensure broader geographical and linguistic representation. The following section will discuss the dataset construction in detail.
@@ 3.1 Operational Definitions
In this study, polarization refers to the process or phenomenon in which opinions, beliefs, or behaviors become more extreme or divided, leading to a greater distance or conflict between differing groups. We also classify the types of polarization as follows:
  • Political/ideological polarization: This type of extremism focuses on division, intolerance, and conflict between political parties and followers.
  • Racial or ethnic polarization: This type of polarization focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races.
  • Religious polarization: This type of polarization focuses on religious identity and incites division, intolerance, and conflict between religious followers
  • Gender polarization: This type of polarization refers to the exclusion, discrimination, and marginalization of individuals based on their gender.
  • Sexual orientation polarization: This refers to the increasing division and distinction between different sexual orientations within society, often leading to heightened tensions, misunderstandings, conflicts, or marginalization among various groups.
  • Other: Polarized texts targeting other groups/identities such as economy class, technology, media, polarization, etc are classified as others.
Furthermore, we define a polarized text manifestations or tactics if it exhibits one or more of the following characteristics:
@@
  • Extreme Language and Absolutism: The use of language that is extreme or makes definitive, all-encompassing statements, often involving words like "always," "never," "worst," or "best," and presenting issues in a dichotomous manner such as "us versus them" or "right versus wrong.
  • Lack of Empathy or Understanding: The absence of compassion or recognition for other viewpoints or experiences in the text.
  • Invalidation: The act of denying or dismissing the identity and existence of individuals or groups, thereby rejecting their sense of self and their presence. See Appendix A for further details and examples for each manifestation of polarization.
@@ 3.2 Data Collection
We collected data from various online platforms, including X (formerly Twitter), Facebook, Reddit, Bluesky, Threads, and news/commentary forums. We use a dynamic keyword-driven strategy tailored for each language. Human experts curated keyword lists to reflect culturally and politically significant discourse across regions and events. Table˜1 illustrates the compositions, size, and event coverage of the dataset.
The Amharic dataset focuses on the Northern Ethiopia or Tigray War in Ethiopia; Arabic texts cover a broad array of social and economic topics; English data centers on US elections and international conflicts; the German corpus features election discourse and migration debates; the Hausa subset captures religious and ideological discussions in West and Central Africa; Spanish content spans abortion, migration, and indigenous/gender rights; and Urdu texts reflect political and sectarian divides.
Language Source(s) Event Train Dev Test Total Instances
Amharic Facebook, X Tigray War in Northern Ethiopia 3500 500 1000 5000
Arabic Facebook, X, Threads, News Social and economic issues 1482 212 424 2118
English X, BlueSky, News US elections and international conflicts 2117 303 605 3025
German X, BlueSky, Reddit Election discourse and migration debates 2426 347 694 3467
Hausa Facebook, X Religious and ideological discussions in Africa 3893 557 1113 5563
Spanish X, BlueSky Abortion, migration, and indigenous/gender rights 1400 201 401 2002
Urdu X Political and sectarian conflicts 1960 280 560 2800
Table 1: Summary of dataset composition: sources, sizes, and event coverage by task and language.
3.3 Annotation Process
Given the cultural and linguistic breadth of POLAR , we developed detailed, multilingual annotation guidelines and deployed a hybrid strategy combining crowdsourcing with trained community annotators. For crowdsourced annotators, we first filtered candidates based on their previous annotation experience and annotation agreement. Specifically, we selected those who had completed prior tasks with an agreement score of at least 0.8. We also conducted a pilot task consisting of 50 items to further filter annotators based on their agreement scores and average labeling time. For community annotators, we provided a training session followed by a pilot task to assess their performance before assigning the next batch of annotation tasks. During the annotation process, we continuously monitored annotator performance in a batch-wise manner to ensure that the overall process remained on track.
Annotation Guidelines: We developed the guidelines in English and Amharic, and then translated and culturally adapted them for each target language. The English annotation guidelines are provided in Appendix A. Annotators were instructed to:
  • Identify whether a text is polarized;
@@ 3.3 Annotation Process
  • If the text is classified as polarized, tag the type of polarization (political, racial/ethnic, religious, gender/sexual identity, other);

### v3

# and Multi-Event Online Polarization

@@ Abstract
Online polarization poses a growing challenge for democratic discourse, yet most computational social science research remains monolingual, culturally narrow, or event-specific. We introduce POLAR, a multilingual, multicultural, and multi-event dataset with over 110K instances in 22 languages drawn from diverse online platforms and real-world events. Polarization is annotated along three axes, namely detection, type, and manifestation, using a variety of annotation platforms adapted to each cultural context. We conduct two main experiments: (1) fine-tuning six pretrained small language models; and (2) evaluating a range of open and closed large language models in few-shot and zero-shot settings. The results show that, while most models perform well in binary polarization detection, they achieve substantially lower performance when predicting polarization types and manifestations. These findings highlight the complex, highly contextual nature of polarization and demonstrate the need for robust, adaptable approaches in NLP and computational social science. All resources will be released to support further research and effective mitigation of digital polarization globally.
POLAR: A Benchmark for Multilingual, Multicultural,
and Multi-Event Online Polarization
Usman Naseem1, Robert Geislinger2, Juan Ren1, Sarah Kohail3, Rudy Garrido Veliz2, P Sam Sahil2,4, Yiran Zhang1, Marco Antonio Stranisci6,7, Idris Abdulmumin5, Özge Alacam8, Cengiz Acartürk9, Aisha Jabr3,Saba Anwar2, Abinew Ali Ayele10, Simona Frenda7,11, Alessandra Teresa Cignarella7, Elena Tutubalina13,14,15, Oleg Rogov13,16,17, Aung Kyaw Htet1, Xintong Wang2, Surendrabikram Thapa18, Kritesh Rauniyar1, Tanmoy Chakraborty19, Arfeen Zeeshan19, Dheeraj Kodati20, Satya Keerthi21, Sahar Moradizeyveh1, Firoj Alam22,23, Arid Hasan24, Syed Ishtiaque Ahmed24, Ye Kyaw Thu25, Shantipriya Parida26, Ihsan Ayyub Qazi27, Lilian Wanzare28, Nelson Odhiambo Onyango28, Clemencia Siro29, Jane Wanjiru Kimani30, Ibrahim Said Ahmad31,32, Adem Chanie Ali2,10, Martin Semmann2, Chris Biemann2, Shamsuddeen Hassan Muhammad33, Seid Muhie Yimam2 1Macquarie University, 2University of Hamburg, 3Zayed University, 4HKBK College of Engineering, 5University of Pretoria, 6University of Turin, 7aequa-tech, 8Bielefeld University, 9Jagiellonian University, 10Bahir Dar University, 11Heriot-Watt University, 12Ghent University, 13AIRI, 14KFU, 15HSE University, 16MTUCI, 17Skoltech, 18Virginia Tech, 19ITT Delhi, 20ABV-IIITM, 21Mahindra University, 22Qatar Computing Research Institute, 23Hamad Bin Khalifa University, 24University of Toronto, 25LU Lab., Myanmar, 26AMD Silo AI, 27Lahore University of Management Sciences, 28Maseno University, 29Centrum Wiskunde & Informatica, 30Jomo Kenyatta University of Agriculture and Technology, 31Bayero University Kano, 32Northeastern University, 33Imperial College London,
@@ 1 Introduction
Online polarization, defined as sharp division and antagonism between social, political, or identity groups, has become a pervasive threat to democratic institutions, civil discourse, and social cohesion worldwide (Waller and Anderson, 2021). It is often fueled by biased or inflammatory content in digital media, strengthening echo chambers and undermining mutual understanding (Garimella, 2018). Polarized discourse amplifies ideological divides and can escalate into hate speech, harassment, and real-world violence Piazza (2023); Martínez-España et al. (2024). Therefore, early detection of polarization is essential for designing interventions that promote healthier online ecosystems.
@@
Figure 1: Pipeline for POLAR construction. Data curation in 22 languages, annotation workflow, and benchmarking.
@@
Despite increasing attention, computational approaches to polarization face several limitations. First, most existing datasets focus on English or high-resource languages, reflecting a widespread trend across NLP tasks that ignores the rich diversity of linguistic and sociocultural contexts in which polarization manifests Simchon et al. (2022); Piazza (2023); Rojo Martínez (2025). Second, previous studies are event-specific or monodomain, such as U.S. elections or Western political debates, limiting their generalizability Demszky et al. (2019); Casal Bértoa and Rama (2021); Sinno et al. (2022); Piazza (2023). Third, the conceptualization of polarization in NLP has largely been binary or topic-focused Hofmann et al. (2022), overlooking the multifaceted ways in which polarization is expressed through vilification, dehumanization, stereotyping, or other rhetorical tactics Donohue and Hamilton (2022). These tactics are often employed in political rhetoric, social debates, or campaigns to solidify support within a group and increase hostility to others.
To address these gaps, we introduce POLAR, a large-scale, multilingual, multicultural, and multievent dataset for fine-grained polarization detection. POLAR supports 22 languages spanning seven language families and balances high-, medium-, and low-resource languages (see Table˜5). The wide extent of our efforts can be seen in Figure 2. Unlike prior work, POLAR supports three complementary tasks:
  • Binary Polarization Detection: Determine whether a given text expresses polarization. We refer to this task as PolarDetect.
@@
  • Polarization Type Classification: Identify the social dimension underlying the polarization (e.g., political, religious, racial). We refer to this task as PolarType.
@@
  • Manifestation Identification: Detect how polarization is rhetorically manifested, including strategies such as stereotyping, deindividuation, vilification, dehumanization, extreme language, and other rhetorical devices. We refer to this task as PolarManifest.
@@
For each task, we develop a cross-cultural annotation protocol tailored for each language’s sociopolitical context. The complete data construction pipeline is illustrated in Figure˜1. We benchmark a range of Small Language Models (SLMs) and Large Language Models (LLMs) under zero-shot and few-shot settings. Our experiments highlight the challenges of generalization and the limitations of current models in capturing nuanced rhetorical patterns across languages. Our contributions are as follows:
@@
  • We release POLAR, the first large-scale, multilingual, fine-grained dataset for polarization analysis across 22 languages and diverse global events, comprising 110K instances.
@@
  • We provide comprehensive benchmarks using state-of-the-art SLMs and LLMs across multiple evaluation settings.
Figure 2: Languages represented in a world map covered by POLAR, covering diverse linguistic and regional contexts. The language and societal context can present itself across varied areas. Language assignments to countries and regions are approximate.
@@ 2 Related Work
Online polarization poses a threat to social cohesion, exacerbated by social media echo chambers and biased content (Waller and Anderson, 2021; Iandoli et al., 2021; Garimella, 2018). As social media and other online platforms become key arenas for political and cultural discourse, the need for early detection and nuanced understanding of polarization has grown significantly. Polarization detection is important for content moderation, peace building, policy development, responsible digital governance and healthy democracy. Foundational research has defined polarization as both intergroup hostility and blind ingroup cohesion (Arora et al., 2022), and has highlighted its relationship with hate speech, fragmentation, and incivility (Mathew et al., 2021).
A growing body of research has documented the role of online spaces in intensifying polarization across different regions (Kubin and von Sikorski, 2021; Barberá, 2020; Gitlin, 2016; Soares and Recuero, 2021). However, most computational work focuses on high-resource languages and event- or region-specific datasets, limiting generalizability (Kubin and von Sikorski, 2021). This leaves a significant gap in our ability to generalize findings across cultures, languages, and events, especially for low-resource languages or multilingual regions.
The lack of standardized datasets across languages has hindered progress in developing and evaluating polarization detection models with cross-lingual or cross-cultural capabilities. Recent shared tasks on hate speech and toxicity (Basile et al., 2019; Pamungkas et al., 2020) have expanded the language and domain coverage, yet remain less fine-grained regarding polarization’s diverse types and rhetorical manifestations. Our work addresses this gap by presenting the comprehensive, fine-grained dataset benchmark for multilingual, multicultural, and multievent online polarization, enabling robust cross-lingual and context-aware modeling.
@@ 3.1 Operational Definitions
In this work, we define polarization as the increasing extremity of opinions, beliefs, or behaviors, resulting in heightened intergroup divisions and conflict. Polarization types are classified as:
  • Political polarization: Focuses on division, intolerance, and conflict between political parties and followers.
  • Racial or ethnic polarization: Focuses on ethnic identity or racial origin and incites division, intolerance, and conflict between ethnic groups or races.
  • Religious polarization: Focuses on religious identity and incites division, intolerance, and conflict between religious followers
  • Gender/ Sexual polarization: Refers to the exclusion, discrimination, and marginalization of individuals based on their gender or sexual orientations.
  • Other: Polarized texts targeting other groups or identities not covered above, such as economic class, technology or media.
In addition to topical categories, we further distinguish polarization by its rhetorical manifestations, defined as follows:
@@
  • Extreme Language and Absolutism: The use of language that is extreme or makes definitive, all-encompassing statements, often involving words like “always”, “never”, “worst”, or “best”, and presenting issues in a dichotomous manner such as “us versus them” or “right versus wrong”.
  • Lack of Empathy: The absence of compassion or recognition for other viewpoints or experiences in the text.
  • Invalidation: The act of denying or dismissing the identity and existence of individuals or groups, thereby rejecting their sense of self and their presence.
Appendix E contains more details and examples for each manifestation of polarization. [^1] Since we define the above polarization manifestations as rhetorical tactics, we have used the terms “manifestation” and “rhetorical tactics” interchangeably.
@@ 3.2 Data Collection
Data Sources: We collected data from a range of online platforms, including major social media sites (e.g., X, Facebook, Reddit, Bluesky, Threads, YouTube comments, Weibo, and Zhihu) and local news or commentary forums (see Table 6 in the Appendix). For several languages (Chinese, Turkish, Polish, Burmese, and Italian), we sampled and re-annotated instances from existing toxic or hate speech datasets, including ToxiCN (Bai et al., 2025), COLD (Deng et al., 2022), the Turkish Hate Speech Dataset (Çöltekin, 2020), Myanmar Hate Speech (Kyaw et al., 2024), HaSpeede2 (Sanguinetti et al., 2020), HODI (Nozza et al., 2023), and BAN-PL (kołos2024banpl).
Event Selection: We curated the dataset to cover diverse real-world events, grounding event selection in the sociopolitical and socioeconomic contexts specific to each language and cultural setting. The data span a broad range of events and issues, including armed conflicts (e.g., the Tigray War in Ethiopia, the Russia–Ukraine conflict, and the Gaza genocide), elections and party politics (e.g., the 2024 U.S. and 2025 German elections), public health crises, large-scale migration, climate change, and broader socioeconomic debates. The dataset also includes discussions related to gender and indigenous rights, religion, and ideology. For some languages, such as Bengali, a broader sampling strategy was adopted due to the lack of sufficiently large event-specific data on the selected platforms.
# Lang. Inner Agr. ($\kappa$) Total PolarDetect PolarType PolarManifest
@@
Polarized (%) Political Racial / Ethnic Religious Polarization Gender / Sexual Other Stereo- type Vilifi- cation Dehuman- ization Extreme Language Lack of Empathy Invalid- ation
1 eng 0.39 4,834 1,767 (37%) 1,726 422 168 108 190 730 1,272 586 1,156 536 879
deu 0.10 4,771 2,274 (48%) 1,959 883 531 281 658 1,728 1,435 712 1,038 1,272 775
urd 0.29 / 0.70 5,346 3,714 (69%) 3,603 2,908 2,954 2,739 2,713 3,328 3,460 2,973 3,324 3,007 3,059
ben 0.59 5,000 2,127 (43%) 1,701 38 97 26 503 298 1,199 535 236 95 89
hin 0.49 4,117 3,510 (85%) 3,051 500 2,417 472 540 2,047 2,683 750 2,082 2,336 2,703
ori 0.46 3,552 1,021 (29%) 744 179 225 119 130 354 385 24 476 56 120
nep 0.79 3,008 1,510 (50%) 518 422 239 158 354 806 947 198 816 318 450
pan 0.55 2,609 1,280 (49%) 803 153 205 291 233 424 1,038 574 624 324 637
ita 0.39 5,038 2,165 (44%) 412 926 368 461 219
spa 0.26 4,958 2,479 (50%) 1,351 945 787 665 665 1,355 1,517 443 1,199 1,187 526
rus 0.39 5,023 1,525 (30%) 696 494 205 284 119
pol 0.46 3,587 1,504 (42%) 1,313 323 131 165 232
fas 0.78 4,943 3,656 (74%) 2,170 120 476 296 1,197 649 2,850 213 835 487 394
2 hau 0.48 5,477 587 (11%) 267 173 139 44 21 234 68 193 165 48 13
arb 0.25 5,070 2,268 (45%) 1,205 874 424 553 847 1,691 1,896 555 1,540 863 411
amh 0.59 4,999 3,747 (75%) 3,339 1,296 99 29 1,239 2,728 2,398 657 1,527 879 799
3 zho 0.64 6,421 3,208 (50%) 376 1,475 127 1,085 552 1,931 1,188 323 522 506 307
mya 0.13 4,334 2,508 (58%) 1,095 228 133 459 1,956
4 swa 0.56 10,487 5,257 (50%) 279 3,721 371 234 833 4,160 4,324 1,340 2,509 3,120 2,456
5 khm 0.83 9,960 9,042 (91%) 1,825 147 336 169 6,565 6,799 152 122 225 1,093 651
6 tel 0.70 3,550 1,885 (53%) 766 603 318 471 842 398 781 88 477 933 809
7 tur 0.46 3,566 1,776 (50%) 1,569 579 557 221 193 1,453 1,169 378 1,575 369 159
Total 110,650 58,810 (53%) 30,768 17,409 11,307 9,330 20,801 31,113 28,762 10,664 20,326 17,429 15,237
Table 1: Number of samples labeled as positive for each annotation task across languages. Inner agreement values denote inter-annotator agreement per language (Fleiss’s $\kappa$ unless otherwise noted). denotes exceptions: German uses Krippendorff’s $\alpha$; Punjabi reports identical Krippendorff’s $\alpha$ and Cohen’s $\kappa$; Urdu reports Fleiss’s $\kappa$ / Cohen’s $\kappa$. Polarization manifestation annotations are not available for Italian, Russian, Burmese, and Polish. Languages are ordered by language families and sub-branches as defined in Table 5.
To collect data for these events, we adopted a dynamic, keyword-driven strategy tailored to each language and topic. Keyword lists were curated by human experts and native speakers to capture culturally and politically salient discourse and were used to retrieve data from online platforms. Table 7 in the Appendix provides an overview of the dataset composition, size, and event coverage.
Dataset Quality Control: To ensure high-quality annotations across languages, we implemented several steps to ensure data quality throughout the quality control process. Before annotation, native speakers developed language-specific preprocessing pipelines. These pipelines included standard NLP procedures such as tokenization, word count filtering, and duplicate detection. Instances that were either too short or excessively long based on language-specific thresholds were removed. For anonymization, all usernames and URLs were replaced with standardized placeholders. For some languages, LLMs were used during the pre-filtering stage to increase the proportion of polarized content.
3.3 Annotation Process and Guidelines
We used a hybrid annotation strategy, leveraging crowd-sourced annotators and trained community annotators for low-resource languages where crowd-sourced annotation support is limited. For the crowd-sourced setting, annotators were selected based on their prior experience and annotation quality. Specifically, we filtered candidates using historical annotation agreement scores and conducted pilot rounds to identify those with consistent performance. Only annotators achieving a Fleiss’ Kappa score of at least 0.8 were retained for the main task. For community annotation, we recruited native speakers with at least a bachelor’s degree. Annotators received training, followed by a pilot round to assess their understanding and performance. Subsequent annotation was assigned in batches, with annotator performance monitored continuously to ensure consistency and adherence to guidelines.
The trained annotators used POTATO Pei et al. (2022) and Label Studio [^2] https://labelstud.io as annotation platforms, while the crowd-sourced annotators used Mechanical Turk [^3] https://www.mturk.com and Prolific [^4] https://www.prolific.com.
Annotation Guidelines: Given the cultural and linguistic breadth of POLAR, we developed detailed, multilingual annotation guidelines (see Appendix E) in English, and then translated and culturally adapted them for each target language. Annotators were instructed to:
  • Identify whether a text is polarized
  • If the text is classified as polarized, tag the type of polarization (political, racial/ethnic, religious, gender/sexual identity, other)